1)The first two terms of a geometric sequence are a1=1/3 and a2=1/6. what is a8 the 8th term?
2)If x is inversely proportional to y and x =60 when y =0.5 find x when x=12
3)An infinite geometric series has 1 ad 1/5 as its first 2 terms: 1, 1/5, 1/25, 1/125,... what is the sum, S, of the infinite series?
4) if y=7.2 when x=10 what is the value of x when y =20? (y varies inversely as the square of x)

Answer by stanbon(57307)

1)The first two terms of a geometric sequence are a1=1/3 and a2=1/6. what is a8 the 8th term?
r=(1/6)/(1/3) = 1/2
a(8)= ar^7 = (1/3)(1/2)^7 = 1/3*1/128= 1/384
----------------
2)If x is inversely proportional to y and x =60 when y =0.5 find x when x=12
y=k/x
0.5=k/60
k=30
So y=30/x
I assume you meant to say "find y when x=12".
If so,
y=30/12 = 5/2
----------------
3)An infinite geometric series has 1 ad 1/5 as its first 2 terms: 1, 1/5, 1/25, 1/125,... what is the sum, S, of the infinite series?
What methods are you using to find the sum of an infinite series?
---------------
4) if y=7.2 when x=10 what is the value of x when y =20? (y varies inversely as the square of x)
y=k/x^2
7.2=k/(100)
k=720
So, y = 720/x^2
20=720/x^2
x^2=720/20
x^2=36
x=6 or x=-6
Cheers
Stan H.